I took a practice test and got 20 WRONG out of 50 QUESTIONS so what would my grade be? Please explain how you got to your answer.
Answer by jim_thompson5910(23738)

If there are 50 questions, and you got 20 wrong, then you got 50-20 = 30 correct.


So the grade as a percentage is (30/50)*100 = 0.6*100 = 60%